### v1

@@ Abstract
Standard formulations of prescribed worst-case disturbance energy-gain control policies for linear time-varying systems depend on all forward model data. In a discrete-time setting, this dependence arises through a backward Riccati recursion. The aim herein is to consider the infinite-horizon $\ell_{2}$ gain performance of state feedback policies with only finite receding-horizon preview of the model parameters. The proposed synthesis of controllers subject to such a constraint leverages the strict contraction of lifted Riccati operators under uniform controllability and observability. The main approximation result establishes a sufficient number of preview steps for the performance loss to remain below any set tolerance, relative to the baseline gain bound of the associated infinite-preview controller. Aspects of the main result are explored in the context of a numerical example.
@@ 1 Introduction
The synthesis of feedback controllers for linear time-varying systems against infinite-horizon quadratic performance criteria has a long history [1, 2, 3, 4, 5]. Generally, the standard formulations are somewhat impractical, as the control policies depend on all future time-varying parameters. In the discrete-time setting of this paper, the stabilizing solution of a backward Riccati recursion encodes this dependence [6, 7, 8, 9].
In periodic settings, including the time-invariant case, availability of the model data across one period suffices to obtain the infinite-horizon solution of the backward recursion via a finite-dimensional algebraic Riccati equation [10, 11, 12, 13]. In more general time-varying settings, an infinite-horizon lifting, given all model data, also leads to algebraic characterization of the solution [14, 15, 16]. However, the practical concern of computation remains a challenge unless a finite structure underlies the parameter variation, e.g., periodicity [16], or more general switching within a finite set [17]. Furthermore, such infinite-horizon lifting would be infeasible in supervisory hierarchies involving online roll-out of finite-horizon plans that determine the model data relevant for low-level control.
This paper concerns state feedback controller synthesis subject to finite receding-horizon preview of the time-varying model data without preview of the disturbance. The goal is to ensure that the closed-loop infinite-horizon energy gain does not exceed a specified worst-case bound. The proposed approach involves approximation of the standard infinite-preview formulation to within desired tolerance, via a Riccati contraction property that holds under uniform controllability and observability and an associated finite-horizon lifting. Related work on the disturbance-free optimal linear quadratic regulator can be found in [18][19, Ch. 4].
@@ 1.1 Basic notation and problem formulation
The following result is the standard infinite-horizon formulation of a (strictly causal) state feedback $\ell_{2}$ gain control policy; e.g., see [4, 5], [9][19, Sec. 2.4.2].
@@ 1.4 Organization
The paper is organized as follows. The main result, outlined in Section 1.2, is formulated as Theorem 2 in Section 2. This section includes development of the underlying finite-horizon lifting used to obtain a one-step controllable and observable model, which enables direct application of an existing result on Riccati operator contraction in the synthesis of the approximating control policy. A structured proof of the main result, based on strict contraction of lifted $\ell_{2}$ gain Riccati operators, is developed in Section 3. A numerical example is considered in Section 4 to explore aspects of the main result, followed by some concluding remarks in Section 5.
@@ 2 Main result
This construction of $X_{t+1}$ involves $d\cdot(T+1)$ steps of model data ahead of $t\in\mathbb{N}_{0}$. The main result formulated below identifies a sufficient number of lifted preview steps $T$ for the corresponding state feedback control policy (10) to meet a specified performance loss bound $\beta\in\mathbb{R}_{>0}$, relative to the infinite-preview policy (8) with prescribed baseline energy-gain performance bound $\gamma$.
@@
Under part 2) of the hypothesis in Theorem 2, $\tilde{X}_{t+1}\in\mathbb{S}_{++}^{n}$ in (33), and $\mathcal{R}_{\gamma,k|t}(X)\in\mathbb{S}_{\smash{++}}^{n}$ over $X\in\mathbb{S}_{\smash{++}}^{n}$, as elaborated in Remark 3.11, Section 3.2. Thus, $X_{t+1}\in\mathbb{S}_{\smash{++}}^{n}$ in (34). Using one-step controllability and observability of the lifted model, it is established in [19, Sec. 5.4.2] that the three components of this part of the hypothesis are necessary conditions when for every $t\in\mathbb{N}_{0}$, the lifted deadbeat policy
@@ 3.1 Lifted Riccati operator contraction
By application of [20, Thm. 1.7], the hypothesis in Theorem 3 is sufficient for strict contraction of the lifted Riccati operator $X\mapsto\tilde{{\mathcal{R}}}_{\gamma,k|t}(X)$ given by (2). Without Assumption 2, the second part of the hypothesis regarding $\tilde{{B}}_{k|t}\tilde{{R}}_{\smash{k|t}}^{-1}\tilde{{B}}_{\smash{k|t}}^{\prime}$ and $\tilde{{Q}}_{k|t}$ is infeasible, and the contraction may be non-strict. The Riemannian distance between $X,P\in\mathbb{S}_{\smash{++}}^{n}$ is defined as follows [36]:
@@ 3.2 Riccati operator lifting
If $\tilde{{R}}_{k|t}\in\mathbb{S}^{(m+n)d}$ is also non-singular, then in (2), $\tilde{{\mathcal{R}}}_{\gamma,k|t}(P)={\mathcal{R}}_{\gamma,k|t}(P)$ by [37, Prop. 12.1.1]. Further,
@@ 5 Conclusion
Application of Weyl’s inequality for singular values [38, Thm. 3.3.2] then yields

### v2

@@ Abstract
Standard formulations of prescribed worst-case disturbance energy-gain control policies for linear time-varying systems depend on all forward model data. In discrete time, this dependence arises through a backward Riccati recursion. This article is about the infinite-horizon $\ell_{2}$ gain performance of state feedback policies with only finite receding-horizon preview of the model parameters. The proposed synthesis of controllers subject to such a constraint leverages the strict contraction of lifted Riccati operators under uniform controllability and observability. The main approximation result is a sufficient number of preview steps for the incurred performance loss to remain below any set tolerance, relative to the baseline gain bound of the associated infinite-preview controller. Aspects of the result are explored in a numerical example.
@@ 1 Introduction
The synthesis of feedback controllers for linear time-varying systems against infinite-horizon quadratic performance criteria has a long history [23, 2, 27, 15, 19]. Generally, the standard formulations are somewhat impractical, as the control policies depend on all future time-varying parameters. In the discrete-time setting of this article, the stabilizing solution of a backward Riccati recursion encodes this dependence [17, 1, 4, 25].
In periodic settings, including the time-invariant case, availability of the model data across one period suffices to obtain the infinite-horizon solution of the backward recursion via a finite-dimensional algebraic Riccati equation [7, 38, 20, 5]. In more general time-varying settings, an infinite-horizon lifting, given all model data, also leads to algebraic characterization of the solution [18, 34, 9]. However, the practical concern of computation remains a challenge unless a finite structure underlies the parameter variation, e.g., periodicity [9], or more general switching within a finite set [10]. Furthermore, infinite-horizon lifting would be infeasible in supervisory hierarchies involving online roll-out of finite-horizon plans that determine the model data relevant for low-level control.
This article concerns state feedback controller synthesis subject to finite receding-horizon preview of the time-varying model data without preview of the disturbance. The goal is to ensure that the closed-loop infinite-horizon energy gain does not exceed a specified worst-case bound. The proposed approach involves approximation of the standard infinite-preview formulation to within desired tolerance, via a Riccati contraction property that holds under uniform controllability and observability and an associated finite-horizon lifting. Related work on the disturbance-free optimal linear quadratic regulator can be found in [36][37, Ch. 4].
@@ 1.1 Basic notation and problem formulation
The following result is the standard infinite-horizon formulation of a (strictly causal) state feedback $\ell_{2}$ gain control policy; e.g., see [15, 19], [25][37, Sec. 2.4.2].
@@ 1.4 Organization
The article is organized as follows. The main result, outlined in Section 1.2, is formulated as Theorem 2 in Section 2. This section includes development of the underlying finite-horizon lifting used to obtain a one-step controllable and observable model, which enables direct application of an existing result on Riccati operator contraction in the synthesis of the approximating control policy. A structured proof of the main result, based on strict contraction of lifted $\ell_{2}$ gain Riccati operators, is developed in Section 3. A numerical example is considered in Section 4 to explore aspects of the main result, followed by some concluding remarks in Section 5.
@@ 2 Main result
This construction of $X_{t+1}$ involves $d\cdot(T+1)$ steps of model data ahead of $t\in\mathbb{N}_{0}$. The main result formulated below identifies a sufficient number of lifted preview steps $T$ for the corresponding state feedback control policy (10) to meet a specified performance loss bound $\beta\in\mathbb{R}_{>0}$, relative to the infinite-preview policy (8) with prescribed baseline energy-gain bound $\gamma$.
@@
Under part 2) of the hypothesis in Theorem 2, $\tilde{X}_{t+1}\in\mathbb{S}_{++}^{n}$ in (33), and $\mathcal{R}_{\gamma,k|t}(X)\in\mathbb{S}_{\smash{++}}^{n}$ over $X\in\mathbb{S}_{\smash{++}}^{n}$, as elaborated in Remark 3.11, Section 3.2. Thus, $X_{t+1}\in\mathbb{S}_{\smash{++}}^{n}$ in (34). Using one-step controllability and observability of the lifted model, it is established in [37, Sec. 5.4.2] that the three components of this part of the hypothesis are necessary conditions when for every $t\in\mathbb{N}_{0}$, the lifted deadbeat policy
@@ 3.1 Lifted Riccati operator contraction
By application of [6, Thm. 1.7], the hypothesis in Theorem 3 is sufficient for strict contraction of the lifted Riccati operator $X\mapsto\tilde{{\mathcal{R}}}_{\gamma,k|t}(X)$ given by (2). Without Assumption 2, the second part of the hypothesis regarding $\tilde{{B}}_{k|t}\tilde{{R}}_{\smash{k|t}}^{-1}\tilde{{B}}_{\smash{k|t}}^{\prime}$ and $\tilde{{Q}}_{k|t}$ is infeasible, and the contraction may be non-strict. The Riemannian distance between $X,P\in\mathbb{S}_{\smash{++}}^{n}$ is defined as follows [3]:
@@ 3.2 Riccati operator lifting
If $\tilde{{R}}_{k|t}\in\mathbb{S}^{(m+n)d}$ is also non-singular, then in (2), $\tilde{{\mathcal{R}}}_{\gamma,k|t}(P)={\mathcal{R}}_{\gamma,k|t}(P)$ by [30, Prop. 12.1.1]. Further,
@@ 5 Conclusion
Application of Weyl’s inequality for singular values [21, Thm. 3.3.2] then yields